OpenStudy (anonymous):

Choose two trinomials from the list below to factor. Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x2 + 13x + 15 2x2 + 27x + 13 5x2 − 2x − 7 10x2 − 7x − 3 8x2 − 11x + 3

OpenStudy (anonymous):

\[2x^2 + 13x + 15 = 2x^2 + 10x +3x + 15 =2x(x+5)+3(x+5) =(2x+3)(x+5)\]

OpenStudy (anonymous):

\[2x^2 + 27x + 13 =2x^2 + 26x +x + 13=2x(x+13)+1(x+13) =(2x+1)(x+13)\]

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

Could u tell me how u got the answer for that owo